OpenStudy (anonymous):

Need to find out how much each box of fruit weighs, and do not know how to start: A+B=48 B+C=54 C+D=46 D+E=32 A+C+E=64

OpenStudy (anonymous):

There are several ways of approaching a large system of equations. Do you know how to reduce matrices into reduced row echelon form?

OpenStudy (mathmate):

You can also solve the system by successive addition and subtraction if you noticed that the variables are chained in a near diagonal form. The answers turned out to be all even (by chance, I guess).

OpenStudy (anonymous):

I really don't know how to even start thinking.... I don't know the rules. For instance...If I take one part of the above problem one step further and say B=48-A and B=C-54. That means that 48-A=C-54 Then I can say C-A=6. My question then is...Could I then draw the conclusion that B=C-A and B=6...? Or am I waaaay out of order?

OpenStudy (callisto):

I'm confused... B=48-A B=C-54 So, 48-A = C - 54 Then, it should be A+C = 48+54, ie, A+C = 102...

OpenStudy (mathmate):

Reduction to the echelon form is the standard solution, perhaps a little more work for a 5x5 system. I will start you off with successive elimination (that does not apply in all cases, only cases similar to this one). A+B =48 ...(1) B+c =54 ... (2) Subtract A -c = -6 (1A) C+D=46 ...(3) Add A +D=40 ...(2a) ... Continue this way until you get A-E=8 ....(3A) add to A+C+E=64 ...(5) to get 2A+C=72 Add to (1A) and solve for A. Substitute A in (1a), (2a), (3a).... to get remaining variables.

OpenStudy (anonymous):

use matrices! they're so much cleaner!

OpenStudy (callisto):

And... wait! B+C = 54 B = 54 - C. not C-54 In that cause B=48-A B=54-C So, 48 - A = 54-C A-C = 48 - 54 A-C = -6 In other words, C-A = 6 , that make sense now :|

OpenStudy (anonymous):

Thank you so much for your responses! Still a bit confused about the rules, but will work on this :0) Thanks again! And Mathmate, thank you so much for the detailed answer, will try to work through that!

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (callisto):

I'm thinking if this way works. A+B=48 -(1) B+C=54 -(2) C+D=46 -(3) D+E=32 -(4) A+C+E=64 -(5) (1)+(2)+(3)+(4) A+E + 2(B+C+D) = 48+54+46+32 (64-C) + 2(B+C+D) = 180 From (2) , B = 54 - C From (3), D = 46 - C So, (64-C) + 2 [(54-C) +C+ (46-C)] = 180 (64-C) + 2 (100-C) = 180 ^ You can solve C from here. Then, solve the equations one by one to get the other values for other unknowns.

OpenStudy (anonymous):

Thank you Callisto! Callisto and Mathmate, I will for sure come back to see you! Now I need to go and work on this! You're the best!